# Algorithmic Collusion And The Minimum Price Markov Game

### v1

@@ 6 Discussion
Nonetheless, it is essential to add nuance to achieve a comprehensive understanding of this complex issue. While legal concerns stem from tangible evidence found in digital open markets and auctions (see Section 1), algorithmic implicit coordination has been shown to be impeded in the specific context of unbiased social dilemmas, as discussed in Section 3. Additionally, the potential for algorithmic and tacit collusion in transparent markets (including public markets) will always exist. Complex cooperative behaviors can be engineered in virtually any large-scale agent system [40], even if such implementations remain challenging. For instance, cooperation in competitive settings can be elusive even when explicit communication is permitted [58, 59]. What is certain, at least from our perspective, is that further research is required to establish the precise theoretical and algorithmic conditions, as well as the practical considerations, regarding algorithmic collusion in public procurement markets.
Extending the MPMG. Future research should aim to extend the MPMG in both quantitative and qualitative dimensions. Quantitatively, this could involve creating environments that more closely mimic actual simulations, using approaches like empirical games [60]. Qualitatively, it could involve relaxing certain assumptions, such as the common value assumption (Assumption 4), particularly in highly heterogeneous market scenarios. Ultimately, developing a realistic, full-scale simulation of such a system would also necessitate a continuous formulation for the action space and a modular environment allowing for the implementation of various supplier selection models, thus generalizing the MPMG to the Public Procurement Markov Game (PPMG).
@@
AI rationality and cyber cartels. In an AI-governed world, the concept of rationality takes on new dimensions, particularly in economic interactions and market behavior. The experiments in this study operate under a set of restrictions to maintain an environment unbiased towards cooperation. However, as mentioned in Section 5, market players could align on algorithmic pricing using centralized learning and execution (i.e., a shared joint action space). This raises the question: could tacit collusion, emerging from a deliberately biased algorithmic structure, be easily detectable?
The potential existence of cyber cartels introduces new research opportunities into how firms might evade detection and punishment by law enforcers through planned algorithmic collusion. Such a consortium could gain advantages in both efficiency and stealth. Explicit coding for collusion would typically require real-life coordination among market players, potentially leaving traces and being susceptible to algorithmic auditing by legal authorities. However, implicit algorithmic coordination can either be programmed or accidental, creating a grey zone that complicates regulatory efforts. It is also worth noting that both digitized and non-digitized markets can be affected, as the only requirement for algorithmic pricing is the availability of data. Distinguishing between intentional and unintentional collusion requires sophisticated analytical tools and a deep understanding of AI behavior to achieve cooperation, highlighting the need for specialized literature and expertise to foster regulatory adaptation.
Competing Interests
The authors declare no competing interests.
Author Contributions
All authors contributed to the study’s conception and design. Computer experiments were designed and performed by Igor Sadoune. The first draft of the manuscript was written by Igor Sadoune and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
Appendix A: Additional Figures
Figure 6.1: Loss values of the dueling DQN during training of the naive D3QN on the 2-player homogenous MPMG (a) and 5-player heterogeneous MPMG (b).
(a) DQN losses on the 2-player homogenous MPMG
(b) DQN losses on the 5-player heterogeneous MPMG
Figure 6.2: Loss values of the dueling DQN during training the D3QN with opponent modeling on the 2-player homogenous MPMG (a) and 5-player heterogeneous MPMG (b).
(a) DQN losses on the 2-player homogenous MPMG
(b) DQN losses on the 5-player heterogeneous MPMG
Figure 6.3: Actor network loss values during training of MAPPO on the 2-player homogenous MPMG (a) and 5-player heterogeneous MPMG (b). Simalarly, graphs (c) and (d) shows the associated critic network loss values for the 2-player homogeneous and 5-player heterogeneous case.
(a) Actor network losses on the 2-player homogenous MPMG
(b) Actor network losses on the 5-player heterogeneous MPMG
(c) Critic network losses on the 2-player homogenous MPMG
(d) Critic network losses on the 5-player heterogeneous MPMG
Appendix B: Implementation Details
In this section, we expose in detail our implementation of the experiments presented in this manuscript. The code associated to this study is written in Python, and can be found at https://github.com/IgorSadoune/minimum-price-game.
E-Greedy Method. Our EpsilonGreedyMAB class utilizes an $\epsilon$-greedy strategy with the following hyperparameters:
@@
Upper Confidence Bound. Our UCBMAB class implements the UCB algorithm with the following hyperparameters:
@@
Thompson Sampling. Our ThompsonSamplingMAB class uses a Bayesian approach to action selection with the following hyperparameters:
@@
Deep Dueling DQN (naive D3QN). The DuelingDQN class implements a Dueling Deep Q-Network with the following architecture:
@@
The D3QNAgent class implements a Double Dueling Deep Q-Network agent with the following hyperparameters:
  • state_dim: Dimensionality of the state space.
  • action_dim: Dimensionality of the action space.
  • lr: Learning rate, set to $1\times 10^{-3}$.
  • gamma: Discount factor, set to 0.99.
  • epsilon: Initial exploration rate, set to 1.0.
  • epsilon_decay: Decay rate of epsilon, set to 0.995.
  • epsilon_min: Minimum exploration rate, set to 0.01.
  • buffer_size: Size of the replay buffer, set to 10,000.
The agent selects actions using an $\epsilon$-greedy strategy:
  • With probability $\epsilon$, select a random action.
  • With probability $1-\epsilon$, select the action with the highest Q-value:
    $a_{t}=\arg\max_{a}Q(s_{t},a)$
The agent stores experiences in a replay buffer and updates the Q-network by sampling mini-batches of experiences. The Q-values are updated using the Bellman equation:
$Q(s_{t},a_{t})=r_{t}+\gamma\max_{a^{\prime}}Q^{\prime}(s_{t+1},a^{\prime})$
where $Q^{\prime}$ is the target Q-network. The loss is computed using Mean Squared Error (MSE):
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left(Q(s_{t}^{i},a_{t}^{i})-y^{i}\right)^{2}$
where $y^{i}=r_{t}^{i}+\gamma\max_{a^{\prime}}Q^{\prime}(s_{t+1}^{i},a^{\prime})$. The target network is updated periodically by copying the weights from the Q-network.
Dueling DQN with opponent modeling (D3QN-OM). Our D3QN-OM shares the same DuelingDQN class with the naive D3QN as MLP representation. However, the OpponentModel class models the behavior of opponents with the following details:
  • state_dim: Dimensionality of the state space.
  • action_dim: Dimensionality of the action space.
  • model: A dictionary that maps states to action distributions.
The opponent model is updated based on observed actions, using the same D3QNAgent class, except for an instance of OpponentModel for opponent modeling. The opponent model is updated based on observed actions as follows:
  1. 1.
    If the current state is not already in the model, initialize the action distribution for this state as an array of zeros.
  2. 2.
    Increment the count for the observed action in the action distribution array corresponding to the current state.
The agent selects actions using an $\epsilon$-greedy strategy, but when opponent modeling is enabled, Q-values are adjusted based on opponent action distributions
$Q(s_{t},a)\times\text{opponent action distribution}$
The D3QN-OM agent interacts with stored experiences in a replay buffer as the same way the naive D3QN does.
Multiagent Proximal Policy Optimization (MAPPO). The MLPNetwork class implements a multi-layer perceptron network with the following architecture:

### v2

@@ 5 Discussion
In this section, we discuss, interpret, and put our results into perspective.
5.1 Interpretation
Three key insights emerge from the computer experiments.
First, the MPMG is a reliable approximation of market dynamics, as it exhibits properties consistent with Jean Tirole’s seminal work on market collusion potential [47]. Specifically, as shown in Section 3.1, in the MPMG, the incentive to defect while opponents cooperate is asymmetrical with respect to the power parameter $\beta$, thereby undermining the intrinsic value of the Pareto strategy. This effect strengthens as the number of agents increases, further reducing the likelihood of successful collusion.
Second, tacit coordination relies heavily on self-reinforcing trends. Indeed, UCB displays counter-intuitive behaviors for a no-regret algorithm. In such settings, the regret of not defecting is higher than the regret associated with not cooperating (see Table 1). Additionally, other no-regret algorithms, TS and $\epsilon$-greedy, favor the Nash Equilibrium. The particularity of UCB, at least in the original version we implemented [54], is that it is a deterministic method because of the way exploration is handled. Therefore, when fully symmetric UCB agents populate the game, coordination towards the Pareto outcome builds upon self-reinforcing beliefs as every agent acts the same. As we can see from Figure 2(a), the confidence intervals for the average (across repeats) joint action frequencies associated with defection and cooperation of the UCB agents are null. In contrast, Figure 2(b) shows non-null confidence intervals for MAPPO agents, and Figure 3 in Appendix A shows that only UCB-related graphs can have null confidence intervals. Furthermore, it has been shown that UCB agents are the RL agents most prone to cooperate in the Iterated Prisoner’s Dilemma, with cooperation rates similar to ours [58].
Figure 2: Average joint action frequencies of UCB (a) and MAPPO (b) agents in the $2$-player homogeneous MPMG.
(a) UCB
(b) MAPPO
Finally, more information does not lead to more cooperation. There appears to be no positive correlation between the sophistication of our AI agents and the emergence of tacit coordination, a conclusion we share with other observers [58, 43, 59]. In fact, in the Prisoner’s Dilemma, no amount of information can a priori defuse the tension created by the existence of the Nash Equilibrium, as long as agents do not engage in self-reinforcing trust over cooperation.
We can therefore conclude that, provided the MPMG reasonably approximates minimum price-ruled auctions, and although collusion can occur in the most naive settings, the minimum price rule is resilient to non-engineered tacit coordination among rational actors, as is any rule preserving the structure of a social dilemma.
5.2 Practical Considerations
This study focuses on evaluating the behavior of artificial agents without explicitly quantifying their collusive potential in terms of collusive gains. Although the Pareto outcome provides higher returns compared to the Nash outcome, agents cannot strictly maximize their returns, as this would involve behavior divergent from typical learning processes. Moreover, even during evaluation—when agents no longer explore and strictly follow their policies—they often converge to mixed strategies rather than a pure Nash or Pareto strategy. Additionally, unlike bandits, D3QN and MAPPO map a policy to state observation, meaning that even when they learn the pure Pareto profile, their collusive gains are highly context-dependent. As a result, their returns in evaluation remain stochastic, and assessing them in terms of collusive gains may be both challenging and irrelevant.
It is also important to note that, in the context of the MPMG, other (non-Nash) suboptimal outcomes can be viewed as rotation schemes, that is, turnovers in the winning set of players. This kind of strategy emerges from mixed behavior and can lead to greater returns than a consistent Nash play.
5.3 Algorithmic Pricing and Legal Perspective
@@
Nonetheless, it is essential to add nuance for a comprehensive understanding of this complex issue. While legal concerns stem from tangible evidence found in digital open markets and auctions (see Section 1), algorithmic implicit coordination has been shown to be impeded in the specific context of unbiased social dilemmas, as discussed in Section 2. Additionally, the potential for algorithmic and tacit collusion in transparent markets (including public markets) will always exist. Complex cooperative behaviors can be engineered in virtually any large-scale agent system [40], even if such implementations remain challenging. For instance, cooperation in competitive settings can be elusive even when explicit communication is permitted [64, 65]. What is certain, at least from our perspective, is that further research is required to establish the precise theoretical and algorithmic conditions, as well as the practical considerations, regarding algorithmic collusion in public
5.4 Future Work
Extending the MPMG.
Future research should aim to extend the MPMG in both quantitative and qualitative dimensions. Quantitatively, this could involve creating environments that more closely mimic actual simulations, using approaches like empirical games [66]. Qualitatively, it could involve relaxing certain assumptions, such as the common value assumption (Assumption 3), particularly in highly heterogeneous market scenarios. Ultimately, developing a realistic, full-scale simulation of such a system would also necessitate a continuous formulation for the action space and a modular environment allowing for the implementation of various supplier selection models, thus generalizing the MPMG to a general auction Markov game.
@@ 5.4 Future Work
AI rationality and cyber cartels.
In an AI-governed world, the concept of rationality takes on new dimensions, particularly in economic interactions and market behavior. The experiments in this study operate under a set of restrictions to maintain an environment unbiased towards cooperation. However, as mentioned in Section LABEL:sec:Computer_Experiments, market players could align on algorithmic pricing using centralized learning and execution (i.e., a shared joint action space). This raises the question: could tacit collusion, emerging from a deliberately biased algorithmic structure, be easily detectable?
The potential existence of cyber cartels introduces new research opportunities into how agents might evade detection and punishment by law enforcers through planned algorithmic collusion. Such a consortium could gain advantages in both efficiency and stealth. Explicit coding for collusion would typically require real-life coordination among market players, potentially leaving traces and being susceptible to algorithmic auditing by legal authorities. However, implicit algorithmic coordination can either be programmed or accidental, creating a grey zone that complicates regulatory efforts. It is also worth noting that both digitized and non-digitized markets can be affected, as the only requirement for algorithmic pricing is the availability of data.
Appendix A Additional Figures
Figure 3: Average (across repeats) joint action frequencies over training episodes. Rows (up to down): UCB, $\epsilon$-greedy, TS, D3QN, MAPPO. Columns (left to right): $(n=2,\sigma(\beta)=0.0)$, $(n=2,\sigma(\beta)=0.5)$, $(n=5,\sigma(\beta)=0.0)$, $(n=5,\sigma(\beta)=0.5)$.
Figure 4: Average (across repeats) training losses. Rows (up to down): D3QN, MAPPO Actor, MAPPO Critic. Columns (left to right): $(n=2,\sigma(\beta)=0.0)$, $(n=2,\sigma(\beta)=0.5)$, $(n=5,\sigma(\beta)=0.0)$, $(n=5,\sigma(\beta)=0.5)$.
Figure 5: Average (across repeats) training policy values. Rows (up to down): D3QN Q-values, MAPPO policy values. Columns (left to right): $(n=2,\sigma(\beta)=0.0)$, $(n=2,\sigma(\beta)=0.5)$, $(n=5,\sigma(\beta)=0.0)$, $(n=5,\sigma(\beta)=0.5)$.
Figure 6: Average (across repeats) training losses. Rows (up to down): D3QN, MAPPO Actor, MAPPO Critic. Columns (left to right): $(n=2,\sigma(\beta)=0.0)$, $(n=2,\sigma(\beta)=0.5)$, $(n=5,\sigma(\beta)=0.0)$, $(n=5,\sigma(\beta)=0.5)$.
Appendix B Implementation Details
To replicate the study, follow GitHub.
E-Greedy Method.
Our EpsilonGreedyMAB class utilizes an $\epsilon$-greedy strategy with the following hyperparameters:
@@ Appendix B Implementation Details
Upper Confidence Bound.
Our UCBMAB class implements the UCB algorithm with the following hyperparameters:
@@
Thompson Sampling.
Our ThompsonSamplingMAB class uses a Bayesian approach to action selection with the following hyperparameters:
@@
Deep Dueling DQN (D3QN).
The DuelingDQN class implements a Dueling Deep Q-Network with the following architecture:
@@
Multiagent Proximal Policy Optimization (MAPPO).
The MLPNetwork class implements a multi-layer perceptron network with the following architecture: